### v2

# Legal Fact Prediction: The Missing Piece in Legal Judgment Prediction

Abstract
Legal judgment prediction (LJP), which enables litigants and their lawyers to forecast judgment outcomes and refine litigation strategies, has emerged as a crucial legal NLP task. Existing studies typically utilize legal facts, i.e., facts that have been established by evidence and determined by the judge, to predict the judgment. However, legal facts are often difficult to obtain in the early stages of litigation, significantly limiting the practical applicability of fact-based LJP. To address this limitation, we propose a novel legal NLP task: legal fact prediction (LFP), which takes the evidence submitted by litigants for trial as input to predict legal facts, thereby empowering fact-based LJP technologies to perform prediction in the absence of ground-truth legal facts. We also propose the first benchmark dataset, LFPBench, for evaluating the LFP task. Our extensive experiments on LFPBench demonstrate the effectiveness of LFP-empowered LJP and highlight promising research directions for LFP. Our code and data are available at https://github.com/HPRCEST/LFPBench.
Legal Fact Prediction: The Missing Piece in Legal Judgment Prediction
1Junkai Liu, 3Yujie Tong, 1Hui Huang, 4Bowen Zheng, 5Yiran Hu, 3Peicheng Wu, 2Chuan Xiao, 2Makoto Onizuka, 1Muyun Yangthanks: Correspondence to: Shuyuan Zheng <zheng@ist.osaka-u.ac.jp>, Muyun Yang <yangmuyun@hit.edu.cn>., 2Shuyuan Zheng11footnotemark: 1 1Harbin Institute of Technology, 2Osaka University, 3Zhejiang University, 4Dalian University of Technology, 5Tsinghua University
Figure 1: Connection between the legal fact prediction (LFP) and legal judgment prediction (LJP) tasks and comparison of three pipelines of LJP: fact-based LJP, evidence-based LJP, and LFP-empowered LJP. Most existing studies focus on fact-based LJP, while evidence-based LJP and LFP-empowered LJP remain unexplored.
1 Introduction
Advancements in NLP technology have significantly propelled the development of legal technology, particularly in the field of legal judgment prediction (LJP). LJP aims to predict court rulings based on litigation case information and legal provisions. For judges, automated predictions can serve as a reference for their official rulings, ensuring consistency in judicial standards. For litigants and their lawyers, pre- or in-trial judgment predictions help assess the potential outcomes of litigation, enabling them to make informed decisions. Consequently, LJP holds great potential for enhancing judicial efficiency and transparency.
Extensive research efforts have been devoted to achieving accurate LJP. However, most of the existing LJP research is limited to fact-based LJP (e.g., (Luo et al., 2017; Zhong et al., 2018; Chen et al., 2019; Yue et al., 2021; Feng et al., 2022; Wu et al., 2022; Gan et al., 2023)), where the input to the LJP system consists of (ground-truth) legal facts, i.e., facts that are formally established through evidence and determined by the judge. However, the users of LJP, such as litigants and their lawyers, typically confirm their legal facts only at a very late stage of the litigation Medvedeva and Mcbride (2023). Consequently, the application of LJP is largely confined, as the users often seek to predict judgments before litigation or in its early stages to develop and adjust litigation strategies or related plans.
To address the limitations of prior LJP studies, this paper proposes a novel legal NLP task: legal fact prediction (LFP), which aims to take the evidence submitted by litigants for trial as input and automatically determine relevant legal facts. Building on this foundation, we further introduce LFP-empowered LJP, as illustrated in Figure 1. In this approach, users first input available evidence into the LFP system to generate predicted legal facts, which are then used as the basis for the subsequent LJP task. This approach aligns more closely with real-world legal practice.
To further facilitate research on LFP and LFP-empowered LJP, this paper introduces the first benchmark dataset, LFPBench, which contains evidence items, legal facts, and judgment outcomes collected from 657 litigation cases in China, covering 10 representative types of civil cases. As such, it can be used to evaluate both the LFP and LJP tasks.
We conducted extensive experiments based on LFPBench, leveraging both general-domain and legal-domain models. The results reveal that, compared to fact-based LJP, evidence-based LJP, where judgments are predicted solely based on evidence, exhibits a significant drop in accuracy. This suggests that the absence of legal facts has a profound impact on LJP. Moreover, LFP-empowered LJP reduces the accuracy drop of evidence-based LJP by 38.5% on average. Therefore, we argue that LFP is a crucial missing piece in the LJP task.
We summarize our contributions as follows:
  • First, we propose a novel task, legal fact prediction (LFP), which can empower LJP applications to operate in a wider range of real-world scenarios.
  • Second, we introduce LFPBench, the first benchmark dataset for studying LFP and LFP-empowered LJP, to support related research.
  • Third, we conduct extensive experiments on LFPBench. Our results confirm the critical role of LFP in the LJP task and reveal the limitations of state-of-the-art (SOTA) models in addressing LFP. These findings offer valuable insights and guidance for future research.
Figure 2: A trial primarily addresses two tasks: determining legal facts and applying the law.
2 The Legal Fact Prediction Task
In this section, we provide background information and formally define the LFP task.
2.1 Background
In this paper, legal fact refers to the facts of the case determined by the judge during litigation, based on the presentation and cross-examination of evidence by the parties in a legal proceeding Korn (1966). In other words, only facts that can be proven by evidence in a court of law can be acknowledged by the judge as legal facts.
As depicted in Figure 2, in civil law countries such as Germany, France, and China, as well as in common law countries like the UK and the US, court proceedings primarily resolve the following two tasks to reach a judgment:
  • Legal fact-finding: Given the evidence presented and the arguments made by both the plaintiff and the defendant, the judge determines the legal facts of the case.
  • Application of law: The judge applies the law to the legal facts to assess the validity of the plaintiff’s claims and make an appropriate judgment.
As evident, legal facts serve as the foundation for the application of law, and before legal fact-finding is complete, it is logically impossible to predict a judgment based on legal facts. In fact, legal facts are usually finalized when the judge reaches a judgment. Therefore, as Medvedeva and Mcbride (2023) has pointed out, utilizing ground-truth legal facts for LJP is impractical. Instead, litigants typically complete evidence collection before litigation or in its early stages, making evidence-based LJP a more reasonable choice. However, existing research on LJP primarily assume the accessibility of legal facts, namely they are all limited to fact-based LJP (e.g., (Luo et al., 2017; Zhong et al., 2018; Chen et al., 2019; Yue et al., 2021; Feng et al., 2022; Wu et al., 2022)), which is mismatched with real-world legal practice.
2.2 Task Definition
Next, we introduce the formal definition of the LFP task. Let $C$ denote the plaintiff’s claims, which determine the scope of the trial and constrain the space of legal facts to be predicted. Let $Z$ denote the list of evidence for trial, which records all available evidence items to be presented and examined in establishing legal facts. Then, the LFP task requires a system $f$ that takes the tuple $(C,Z)$ as input and yields a set of legal facts $f(C,Z)$.
The predicted legal facts $f(C,Z)$, along with the claims $C$, can be further input into a given LJP system $g$ to obtain the judgments $g(C,f(C,Z))$ for the claims. Since our motivation is to enhance evidence-based LJP, the objective of the LFP task is to find the optimal LFP system $f$ that maximizes the accuracy of the predicted judgments $g(C,f(C,Z))$.
Note that the LFP task is not to summarize the evidence into legal facts. Instead, the available evidence items represent fragmented pieces of legal facts rather than a complete picture. Therefore, the LFP task involves deducing and expanding the evidence into legal facts. Moreover, conflicts or contradictions may exist among evidence items, particularly between the evidence presented by the plaintiff and the defendant. This requires the LFP system to assess the strength and logical coherence of the evidence to resolve the conflicts, making the prediction of legal facts a significant challenge.
Notice the plaintiff’s claims and the evidence list are typically available before the trial, therefore leveraging this information as input for LFP aligns with real-world legal practice.
3 LFPBench
Table 1: Data statistics of the LFPBench dataset. Complete-win: all claims of the plaintiff are supported. Partial-win: part of the plaintiff’s claims are completely or partially supported. Loss: all claims of the plaintiff are rejected.
No. of cases No. of evid. items (plaintiff)
Total 657 Max 19
With Defendant Evid. 387 Avg. 4.26
With Third-Party Evid. 80 Median 4
No. of cases results No. of evid. items (defendant)
Complete-win cases 166 Max 14
Partial-win cases 397 Avg. 1.83
Loss cases 94 Median 1
No. of judgment No. of claims
Full support 631 Max 9
Partial support 621 Avg. 2.48
Reject 378 Median 2
Figure 3: Case type distribution of the LFPBench dataset.
Figure 4: A data sample from the LFPBench dataset featuring a house lease case. Both the plaintiff and the defendant submitted evidence to assist the judge in determining the legal facts. However, Evidence 4, Evidence 6, and Evidence 7 present conflicting information regarding the defendant’s actual move-out date. Ultimately, according to Evidence 7, the judge determined that the defendant had not moved out before July 7 and had defaulted on the rent for June.
3.1 Dataset Overview
Existing LJP datasets, such as CAIL2018 Xiao et al. (2018) and ILDC CJPE Malik et al. (2021), do not include the evidence items submitted by litigants, making LFP infeasible. Therefore, we propose the first benchmark dataset for the LFP task, LFPBench. LFPBench consists of data for 657 first-instance cases in China, covering ten types of civil causes of action as shown in Figure 3. Each case includes the plaintiff’s claims, the evidence items submitted by the litigants, ground-truth legal facts, ground-truth judgments for the claims, and more. Therefore, LFPBench can be used for evaluating both LFP and LJP tasks.
The data statistics of LFPBench can be found in Table 1. In LFPBench, defendants present counterevidence in 58.9% of cases, leading to disputes over the determination of legal facts. Consequently, only 38.71% of claims are fully supported, and plaintiffs completely win only 25.27% of the cases. Therefore, predicting the legal facts and judgments of these cases is highly challenging. We also present a data sample of a house lease case in Figure 4, illustrating how conflicting evidence between the litigants complicates the determination of legal facts.
3.2 Dataset Construction
LFPBench data was extracted from judicial judgments in China. Specifically, we selected ten representative civil litigation causes of action and retrieved 100 written judgments for each type from the China Judgments Online database [^1] https://wenshu.court.gov.cn/. These case types encompass various common disputes over property and personal rights in daily life. To ensure quality, three legal experts reviewed the judgments and excluded those with overly vague descriptions of evidence. Ultimately, 657 cases were retained.
Then, we used regular expressions to extract legal facts and judgment outcomes, as they are typically written in a consistent structure in the documents. Conversely, since the writing format for claims and evidence varies across judgments, we employed GPT-4o to extract this information. Afterward, we conducted a manual review to ensure data integrity.
Finally, our legal experts annotated the judgment outcomes. Specifically, as shown in Figure 4, they aligned each claim with its corresponding judgment and categorized the outcome into three labels based on the level of support: fully supported, partially supported, and rejected. These labels enable us to evaluate the LFP task using classification-based assessment methods. More details are provided in Appendix A.
Table 2: Accuracy (%) of predicted judgments under different LJP approaches and models. Def.: cases where both parties have submitted evidence. No def.: cases where the defendant has not submitted evidence.
Model Evidence-based LFP-empowered Fact-based
All Def. No def. All Def. No def. All Def. No def.
GPT-4o 50.67 50.91 50.31 51.47 49.39 54.67 55.77 52.02 61.53
Claude3.5 50.80 46.36 57.63 52.58 48.58 58.72 56.44 51.21 64.49
Qwen2.5-14B 45.09 42.21 49.53 48.10 43.83 54.67 49.45 44.74 56.70
Llama3.1-Chinese-8B 40.31 34.72 48.91 40.49 34.62 49.53 40.18 35.53 47.35
Average (General) 46.72 43.55 51.60 48.16 44.11 54.40 50.46 45.88 57.52
Law-Llama3.1-8B 31.10 28.85 35.36 30.12 26.72 40.65 33.13 32.89 33.49
LawJustice-Llama3.1-8B 35.21 30.97 41.74 28.96 26.42 32.87 32.33 26.42 41.43
Average (Legal) 33.16 29.91 38.55 29.54 26.57 36.76 32.73 29.66 37.46
(a) ChatLaw embeddings.
(b) OpenAI embeddings.
Figure 5: The correlation between the similarity between predicted and ground-truth legal facts and the accuracy of judgment prediction. The embeddings are derived from two different models to enhance credibility.
4 Experiment
4.1 Setup
Research questions.
We conduct experiments to answer the following questions.
  • RQ1 (Model & LJP Approach Comparison): How do SOTA models perform in LFP? How do different LJP approaches (evidence-based LJP, fact-based LJP, and LFP-empowered LJP) perform?
  • RQ2 (Challenge & Bias Analysis): What are the challenges of the LFP task? What biases do existing models exhibit when handling the LFP task?
Models.
We employ 6 LLMs as the LFP and LJP systems, including the closed-source, general-purpose LLMs GPT-4o-0513 Hurst et al. (2024) and Claude-3.5-Sonnet-20241022222https://docs.anthropic.com/en/docs/about-claude/models, the open-source, general-purpose LLMs Qwen2.5-14B-Instruct Yang et al. (2024) and Llama3.1-Chinese-8B UnicomAI (2024), as well as the open-source legal LLMs LawJustice-Llama3.1-8B BAAI (2024b) and Law-Llama3.1-8B basuo (2024). Additionally, we evaluated other legal LLMs including DISC-LawLLM Yue et al. (2024), Lawyer-Llama-13B-V2 Huang et al. (2023) and AIE-51-8-Law-Model lingminai (2025). However, these models failed to perform the LFP task due to poor instruction-following capabilities, as detailed in Appendix B.2.
LJP approaches.
We compare the following LJP approaches that differ in their input, with their prompts detailed in Appendix B.3.
  • Evidence-based LJP: The submitted evidence and the plaintiff’s claims are directly input into the LJP system to generate legal judgments.
  • LFP-empowered LJP: The submitted evidence and the plaintiff’s claims are first input into the LFP system to predict legal facts. The predicted facts, along with the original inputs, are then fed into the LJP system to generate legal judgments.
  • Fact-based LJP: The LJP system predicts legal judgments based on the ground-truth legal facts and the plaintiff’s claims.
Table 3: Accuracy (%) of LFP-empowered LJP for different judicial cases. Def.: cases where both parties have submitted evidence. No def.: cases where the defendant has not submitted evidence.
Model Complete-win case Partial-win case Loss case
All Def. No def. All Def. No def. All Def. No def.
GPT-4o 66.95 31.06 88.29 44.49 47.85 38.02 60.47 71.52 29.82
Claude3.5 84.75 68.94 94.14 43.83 44.13 43.25 42.79 51.27 19.30
Qwen2.5-14B 75.99 60.61 85.14 45.15 45.99 43.53 16.74 20.25 7.02
Llama-3.1 75.42 59.09 85.14 34.31 34.38 34.16 13.49 15.19 8.77
Law-Llama3.1-8B 51.69 40.15 58.56 31.20 29.23 34.99 5.12 4.43 7.02
LawJustice-Llama3.1-8B 41.81 40.15 42.79 30.16 29.23 31.96 1.86 2.53 0.00
Average 66.10 50.00 75.68 38.19 38.47 37.65 23.41 27.53 11.99
Table 4: Accuracy (%) of LFP-empowered LJP for different causes of action. LPR: Labor Payment Recovery. PC: Pre-sale Contract. SC: Sales Contract. ID: Inheritance. HL: House Lease. TL: Tort Liability. UE: Unjust Enrichment. PR: Property Return. MP: Marital Property. RLBH: Right to Life/Body/Health.
Model LPR PC SC ID HL TL UE PR MP RLBH
GPT-4o 60.16 52.10 62.84 50.58 50.00 54.60 53.59 53.23 39.88 40.40
Claude3.5 66.41 61.08 63.51 55.81 54.49 45.98 51.63 51.08 38.15 41.72
Qwen2.5-14B 71.88 53.89 63.51 47.09 56.18 40.80 41.18 41.94 39.88 30.46
Llama3.1-Chinese-8B 63.28 45.51 52.03 45.35 39.33 34.48 32.03 33.87 31.21 34.44
Law-Llama3.1-8B 47.66 22.75 41.22 31.98 30.34 33.91 29.41 24.73 33.53 31.79
LawJustice-Llama3.1-8B 42.19 26.95 35.81 26.74 24.72 29.31 24.18 22.58 34.10 27.15
Average 58.60 56.37 53.15 42.93 42.51 39.85 38.67 37.91 36.13 34.33
Table 5: The proportions of cases in which each model predicts a complete win, a partial win, or a loss for the plaintiff using LFP-empowered LJP. Evidence items are ordered based on their sources, ensuring that all evidence from the plaintiff (or defendant) appears first.
Model Defendant first, plaintiff last Plaintiff first, defendant last
Accuracy
Complete
win rate
Loss rate
Partial
win rate
Accuracy
Complete
win rate
Loss rate
Partial
win rate
GPT-4o 42.19 16.26 29.75 53.99 48.80 12.27 28.83 58.90
Claude3.5 39.78 42.33 1.84 55.83 49.64 37.42 18.40 44.17
Qwen2.5-14B 45.55 29.14 1.23 69.63 44.35 26.69 2.15 71.17
Llama3.1-Chinese-8B 35.23 19.69 2.77 77.54 34.74 22.09 2.45 75.46
Law-Llama3.1-8B 28.61 4.45 0.40 95.14 27.04 4.60 1.15 94.25
LawJustice-Llama3.1-8B 25.36 3.07 0.00 96.93 28.49 2.03 0.0 97.97
Average 36.12 19.16 5.99 74.51 38.84 17.52 8.83 73.65
Ground truth 100.00 17.18 17.18 65.64 100.00 17.18 17.18 65.64
(a) Bias in complete-win rate.
(b) Bias in partial-win rate.
(c) Bias in loss rate.
Figure 6: The effect of the difference in the number of evidence items between the plaintiff and the defendant on the judgments yielded by LFP-empowered LJP. The bias in y-axis means the difference in the rate between the model’s predictions and the ground truth.
4.2 Model & LJP Approach Comparison
Finding 1: Legal LLMs perform poorly in LFP and LFP-empowered LJP.
As shown in Table 2, the closed-source models GPT-4o and Claude3.5 consistently achieve the best performance across different LJP approaches, while the open-source legal LLMs Law-Llama3.1-8B and LawJustice-Llama3.1-8B perform the worst, with accuracy close to random guessing. This discrepancy cannot be attributed solely to the relatively small size of the legal LLMs, as Llama3.1-Chinese-8B, which has the same model size, performs significantly better. One possible explanation is that these legal LLMs are typically fine-tuned on short-text legal QA datasets BAAI (2024a); Yue et al. (2024); Huang et al. (2023), making them less capable of handling complex tasks such as LFP and LJP, which require summarization, reasoning, and deduction over long texts. Future research could explore incorporating general-domain instruction data into fine-tuning to mitigate the catastrophic forgetting of fundamental capabilities. It is also promising to develop more complex and higher-quality reasoning datasets in the legal domain. This would enable legal models to move beyond simple QA tasks and instead leverage their enhanced legal knowledge to analyze problems step by step, ultimately evolving into domain experts that outperform general models.
Finding 2: Incorporating LFP can substantially reduce the performance gap between evidence-based LJP and fact-based LJP.
As shown in Table 2, for the general-purpose LLMs, predicting legal judgments directly from evidence results in a $7.42\%$ decrease in accuracy compared to predictions based on ground-truth legal facts. This indicates that while fact-based LJP research has made considerable progress, its effectiveness heavily relies on the accessibility of legal facts. On the other hand, although the more practice-aligned evidence-based LJP underperforms [^3] Note that for the legal LLMs, evidence-based LJP outperforms the other approaches. However, given their poor performance close to random guessing, this difference is likely due to randomness rather than a meaningful advantage., the incorporation of LFP reduces the performance gap by 38.50%. Therefore, predicting legal facts from evidence first and then making legal judgments based on the predicted facts can significantly improve the accuracy of evidence-based LJP. Compared to fact-based LJP, LFP-empowered LJP accommodates a broader range of LJP scenarios in legal practice, striking a favorable balance between accuracy and applicability.
Finding 3: More accurate legal facts yield more accurate legal judgments.
We convert both the predicted legal facts and the ground truth into embeddings based on ChatLaw-Text2Vec Cui et al. (2023) and OpenAI text-embedding-3-large OpenAI (2024), and then compute their cosine similarity. Figure 5 shows a positive correlation between the similarity of legal facts and the accuracy of legal judgments. This confirms that high-quality predicted legal facts serve as a crucial foundation for precise LJP.
4.3 Challenge & Bias Analysis
Finding 4: Judgement prediction for partial-win or loss cases is more challenging.
In Table 3, we report the performance of the LLMs in LFP-empowered LJP across cases where the plaintiff achieves a complete win, a partial win, or a loss. The results show that accuracy is typically highest in complete-win cases, while loss cases are the most challenging. This may be because, in real-world scenarios, complete-win cases usually have strong supporting evidence, making it easier to infer legal facts and judgments. In contrast, partial-win and loss cases often involve significant disputes between the parties, with conflicting or contradictory evidence, making it difficult to establish legal facts and reach a judgment.
Finding 5: Judgement prediction for cases with defendant evidence is more challenging.
In Tables 2 and 3, we distinguish between cases where the defendant did or did not submit evidence. The results show that LFP-empowered LJP performs better in the former. This further suggests that counterevidence presented by the defendant can hinder LLMs’ ability to infer legal facts. Therefore, future research should focus on enhancing LLMs’ reasoning capabilities to better assess the authenticity of information.
Finding 6: Judgment prediction with weaker evidence is more challenging.
From Table 4, we observe that cases involving LPR, PC, SC, ID, and HL, which typically feature strong written evidence (e.g., contracts and wills), allow for easier prediction of legal facts and judgments. Moreover, due to their textual nature, written evidence is more easily understood by LLMs. In contrast, cases such as TL and RLBH, which involve torts, often rely on non-written evidence, such as physical objects and audiovisual materials. When described in text, these forms of evidence lose significant detail. This suggests that future research could leverage multimodal technology to better interpret image- and sound-based evidence.
Finding 7: Bias arises from the presentation order of evidence items.
In Table 5, we select all cases that include both plaintiff and defendant evidence, sort the evidence list in different orders, and then feed them to the models for LFP-empowered LJP. We find that the order in which plaintiff and defendant evidence appears significantly influences the predicted judgments, introducing a bias favoring the party whose evidence is presented last. Specifically, when plaintiff evidence appears later, the LLMs are more likely to predict a complete or partial win for the plaintiff. Conversely, when defendant evidence appears later, the likelihood of the plaintiff losing the case increases. This bias may be attributed to the attention mechanism of LLMs Yu et al. (2024), which requires further exploration in future research.
Finding 8: Bias arises from the number of evidence items.
Figure 6 illustrates the impact of the difference in the number of evidence items between the plaintiff and the defendant on the predicted judgments. We observe that as the gap in evidence quantity increases, the LLMs generally tend to predict judgments with a higher complete-win rate and a lower partial-win rate compared to the ground truth. This suggests that an advantage in evidence quantity may lead LLMs to develop a bias toward fully supporting the plaintiff. However, the effect of this advantage on the loss rate is highly divergent: as the plaintiff’s advantage increases, the two closed-source models become less likely to predict a loss for the plaintiff, whereas the open-source models exhibit the opposite trend. Regardless of the model type, further efforts are needed to eliminate these biases and ensure that LLMs understand that an advantage in the number of evidence items does not necessarily indicate stronger evidence, nor does it have any causal relationship with the legal facts.
5 Discussion
Although the input for LFP is defined as the evidence list and the plaintiff’s claims, other trial-related information could also be incorporated into this task. As discussed by Medvedeva and Mcbride (2023), the ideal input for LJP should encompass any information available to the court or the parties at the time of performing LJP, such as complaints, defenses, and evidence submitted by the parties. This principle also applies to LFP. However, the information available to the court or the parties depends on the stage of the trial. For example, before filing a lawsuit, the plaintiff and the defendant may only have access to the evidence they personally possess. After filing, they gain access to each other’s evidence and arguments regarding the legal facts. In this work, we chose the evidence list as the basic input for LFP because, at different stages of the trial, both parties have access to certain evidence.
Note that the "evidence list" here does not necessarily correspond to the final list of evidence submitted to the court, but rather represents the set of evidence items available to the parties at the current stage. Additionally, if the parties have access to other trial-related information, it can be incorporated as supplementary input to improve prediction accuracy. This suggests that in future work, we can adapt the LFP task to different trial stages by tailoring the input, thereby addressing various demands in legal practice.
6 Related Work
Legal Fact Prediction
Research on LJP can be traced back to the 1960s (Lawlor, 1963), which is one of the most fundamental tasks in legal AI. As judgment documents have become publicly accessible in many countries, researchers have extracted legal facts and judgment results from these documents, forming plenty of benchmark datasets for LJP research (e.g., (Xiao et al., 2018; Chalkidis et al., 2019; Malik et al., 2021; Semo et al., 2022; Chalkidis et al., 2022; Hwang et al., 2022)). Using judgment documents, numerous legal NLP studies have explored fact-based LJP, which predicts legal judgments based on legal facts, achieving promising predictive accuracy (e.g., (Luo et al., 2017; Zhong et al., 2018; Chen et al., 2019; Yue et al., 2021; Feng et al., 2022; Wu et al., 2022; Gan et al., 2023)). However, legal facts are not objective facts and are often difficult for LJP’s intended users to obtain before a judgment is rendered. Consequently, Medvedeva et al. (Medvedeva and Mcbride, 2023) recently pointed out that most existing LJP studies rely on unrealistic input such as legal facts, limiting their practicality. Several studies employed legal briefs (Tippett et al., 2021), complaint documents (McConnell et al., 2021), court debate records (Ma et al., 2021) for LJP, but these types of judicial documents are typically not publicly accessible, making it impractical to obtain large-scale datasets for training LJP models. To address the lack of practicality in current LJP research, this paper proposes the LFP task as a preliminary step to fact-based LJP. LFP-empowered LJP establishes a practical loop from evidence to judgment, making LJP more applicable in real-world scenarios.
Legal Document Summarization
Legal Document Summarization (LDS) is the most correlated task with our LFP, which aims to automatically producing concise, accurate, and coherent summaries of legal texts (Kanapala et al., 2019), such as judgment documents Polsley et al. (2016), contracts Manor and Li (2019), and court debate records Duan et al. (2019). Intuitively, while LFP enriches and assembles fragmented and concise pieces of evidence into complete legal facts, LDS takes the opposite approach by refining legal texts and eliminating lengthy and complex details. Additionally, LFP requires inferring logically coherent factual information from conflicting or contradictory evidence, making it more challenging than text summarization.
7 Conclusion
This paper introduces the LFP task to automate the prediction of legal facts for the subsequent LJP task, addressing recent concerns that using ground-truth legal facts for LJP is impractical. We constructed a benchmark dataset for the LFP task, LFPBench, based on publicly available judicial documents. Extensive experiments conducted on LFPBench reveal that SOTA LLMs struggle to accurately determine legal facts when faced with conflicting or contradictory evidence. Moreover, these models exhibit biases related to the quantity and presentation order of evidence. Future work includes addressing the above limitations and constructing larger-scale LFP datasets to facilitate more extensive research.
Limitations
As the first step on the LFP task, this work has the following limitations. First, the evidence information in LFPBench is extracted from publicly accessible judgment documents in China, which are typically summarized and may lack many details of the original evidence, making it more challenging to predict legal facts. Second, the scope of case types covered by LFPBench remains limited, as it does not include criminal or administrative litigation cases.
Ethical Considerations
Our dataset is constructed based on judicial documents, which may present two ethical considerations: (1) Data Privacy and Confidentiality: Judicial documents often contain some basic personal information of litigants, such as names, addresses, and identity numbers. Although we have processed the data to remove or anonymize personally identifiable information (PII), we must still comply with data usage regulations and refrain from any de-anonymization attempts that could compromise personal privacy. (2) Bias and Fairness: Judicial datasets may inadvertently perpetuate biases present in historical legal decisions, related to gender, socioeconomic status, or other factors. Regular audits and updates can be conducted to minimize the risk of biased content.
Appendix A Additional Details on LFPBench
A.1 Annotation
Our dataset’s annotators consist of four graduate students, two of whom have academic background in law, and two in computer science. They are all co-authors of this thesis, and therefore, no remuneration was provided. Before beginning the annotation process, the annotators unified the criteria for "supported", "partially supported", and "rejected". Generally speaking, if there is any conflict between the court’s ruling and the content of a claim (such as a minor discrepancy in the amount of money), it cannot be considered that the court supports the claim. Correspondingly, if there is any overlap between the court’s ruling and the content of a claim (such as the recognition of a small portion of the damages), it cannot be considered that the court rejects the claim.
A.2 Anonymization
Before extracting the relevant legal judgment information, we have already removed and replaced all sensitive information, such as the names and identification numbers of the litigants. Therefore, using this benchmark is safe and does not pose any risk of personal information leakage.
A.3 Copyright Issue
The case data used in LFPBench is sourced from the China Judgments Online (https://wenshu.court.gov.cn/), a website established by the Supreme People’s Court of China for the publication of judgments issued by courts at all levels in China. According to Article 5 of the Copyright Law of the People’s Republic of China: "This Law does not apply to: (1) laws, regulations, resolutions, decisions, orders, and other documents of a legislative, administrative, or judicial nature, and their official translations…" As such, the judgment data used in this paper falls under the exemption outlined in this provision and is not subject to the Copyright Law.
Appendix B Additional Details on Experiments
B.1 Documentation of LLMs
Table 6: Documentation of the used LLMs.
Artifacts License Parameter Scale Language
Lawyer-Llama-13B-V2 Apache License 2.0 13B Chinese
DISC-LawLLM Apache License 2.0 13B Chinese
Llama3.1-Chinese-8B Apache License 2.0 8B Chinese
LawJustice-Llama3.1-8B Apache License 2.0 8B Chinese/English
Law-Llama3.1-8B Not specified 8B Chinese
AIE-51-8-Law-Model Not specified 3B Chinese
Qwen2.5-14B Apache License 2.0 14B Multilingual
GPT-4o Closed-source Unknown Multilingual
Claude3.5 Closed-source Unknown Multilingual
Table 6 shows the basic information of the LLMs involved in our paper. The Apache License 2.0 allows us to freely use these assets for academic research. One of the goals in training domain-specific models is to achieve performance comparable to larger models within specific domains using fewer parameters. Therefore, we have opted for a model with a relatively small parameter count, ranging between 8-14B. All these models possess Chinese language capabilities, making them suitable for our benchmark tests. We report the hyperparameters used for the LLMs in Table 7.
Table 7: Hyperparameter settings.
parameter Close-source Models Open-source Models
frequency_penalty 0 0
logprobs false false
presence_penalty 0 0
temperature 1 0.7
max_output_tokens 4,096 2048
top_p 1 0.8
B.2 Selection of Legal LLMs
We conducted a preliminary evaluation to select legal LLMs. Each model was tested on the entire LFPBench dataset to evaluate its performance on the LJP task, and we calculated the percentage of its outputs that could be successfully extracted by the evaluation script. If the model’s results were recognized by the script, it indicates that the model could follow our instructions for LJP.
Table 8 reports the performance of the general-purpose LLM Llama3.1-Chinese-8B and five legal LLMs in the preliminary evaluation, including DISC-LawLLM Yue et al. (2024) and Lawyer-Llama-13B-V2 Huang et al. (2023), Law-Llama3.1-8B, LawJustice-Llama3.1-8B, and AIE-51-8-Law-Model lingminai (2025). The results show that while Llama3.1-Chinese-8B could fully follow our instructions, the long and complex instructions posed significant challenges for the legal LLMs. Finally, we selected the two legal LLMs with the highest success rate in instruction following for the main experiments in Section 4.
Table 8: Success rate (%) of different legal domain models in following our LJP instructions. For reference, the general-purpose open-source model Llama3.1-Chinese-8B fully complies with our instructions.
Models Base Models Instruction Following Rate (%)
Llama3.1-Chinese-8B Llama3.1-8B (2024) 100.00
Law-Llama3.1-8B Llama3.1-8B (2024) 88.74
LawJustice-Llama3.1-8B Llama3.1-8B (2024) 80.06
DISC-LawLLM Baichuan-13B (2023) 70.47
Lawyer-Llama-13B-V2 Llama2-13B (2023) 62.42
AIE-51-8-Law-Model Qwen2.5-3B (2024) 51.09
B.3 Prompt Templates
In this section, we present the prompt templates we used for evidence extraction, claim extraction, legal fact prediction, evidence-based LJP and fact-based LJP. We have performed initial adjustment to the prompt templates to ensure the performance of different models.
Prompt 1: Evidence Extraction [Court Record] **the original text of the reference judgment paper** [Evidence List] **the reference evidence list** Please follow the format of the example above to extract a list of evidence from the provided trial records and output it in the form of a JSON list. Each element in the list should be a dictionary representing a piece of evidence, containing two key-value pairs: "Party Submitting Evidence" and "Content of Evidence." The Party Submitting Evidence should be one of [Plaintiff, Defendant, Third Party], while the Content of Evidence should be extracted directly from the original text of the trial records. [Court Record] **the original text of the target judgment paper** [Evidence List]
Prompt 2: Claim Extraction [Court Record] **the original text of the reference judgment paper** You need to extract all claims of plaintiff from the court record given above, and then organize them into such a list: [Plaintiff’s Claims] [ "**claim1**", "**claim2**", ] Each claim in the list should be as faithful to the original text as possible. Focus on the subjective opinions put forward by the defendant, and do not pay attention to his specific evidence. You only need to output the formatted list of claims, without adding any comments. [Plaintiff’s Claims]
Prompt 3: Legal Fact Prediction [Plaintiff’s Claims] (1) **claim1** (2) **claim2** [Litigant] **the parties concerned** [Evidence List] (1) **submitting party** **content** (2) **submitting party** **content** Please analyze the plaintiff’s claims and the list of evidence in the above case, and output a faithful description of the basic facts of the case from the court’s perspective. Only provide the findings of fact, without adding any reasoning process or explanations.
Prompt 4: Evidence-Based LJP [Case Type] **one of the ten types** [Litigant] **the parties concerned** [Evidence List] (1) **submitting party** **content** (2) **submitting party** **content** [Plaintiff’s Claims] (1) **claim1** (2) **claim2** You need to refer to the evidence presented by all parties in the [Evidence List] to predict the court’s judgment on the [Plaintiff’s Claims], and form a corresponding judgment list. The [Judgment List] is a list composed of three numbers (0, 1, -1). If you believe the court will fully support the claim, the result is 1; if partially supported, the result is 0; if the claim is dismissed, fill in -1. Just output the formatted judgment list without any comments.
Prompt 5: Fact-Based LJP [Case Type] **one of the ten types** [Litigant] **the parties concerned** [Reference Facts] **the fact determined by court** [Plaintiff’s Claims] (1) **claim1** (2) **claim2** You need to refer to the evidence presented by all parties in the [Evidence List] and the reference facts provided by the model in the [Reference Facts] to predict the court’s judgment on the [Plaintiff’s Claims] and form a corresponding judgment list. The [Judgment List] is a list composed of three numbers (0, 1, -1). If you believe the court will fully support the claim, the result is 1; if partially supported, the result is 0; if the claim is dismissed, fill in -1. Just output the formatted judgment list without any comments.